# Dual-Balancing for Physics-Informed Neural Networks

### v1

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as a new learning paradigm for solving partial differential equations (PDEs) by enforcing the constraints of physical equations, boundary conditions (BCs), and initial conditions (ICs) into the loss function. Despite their successes, vanilla PINNs still suffer from poor accuracy and slow convergence due to the intractable multi-objective optimization issue. In this paper, we propose a novel Dual-Balanced PINN (DB-PINN), which dynamically adjusts loss weights by integrating inter-balancing and intra-balancing to alleviate two imbalance issues in PINNs. Inter-balancing aims to mitigate the gradient imbalance between PDE residual loss and condition-fitting losses by determining an aggregated weight that offsets their gradient distribution discrepancies. Intra-balancing acts on condition-fitting losses to tackle the imbalance in fitting difficulty across diverse conditions. By evaluating the fitting difficulty based on the loss records, intra-balancing can allocate the aggregated weight proportionally to each condition loss according to its fitting difficulty levels. We further introduce a robust weight update strategy to prevent abrupt spikes and arithmetic overflow in instantaneous weight values caused by large loss variances, enabling smooth weight updating and stable training. Extensive experiments demonstrate that DB-PINN achieves significantly superior performance than those popular gradient-based weighting methods in terms of convergence speed and prediction accuracy. Our code and supplementary material are available at https://github.com/chenhong-zhou/DualBalanced-PINNs.
@@ 1 Introduction
Accurately and efficiently producing solutions to partial differential equations (PDEs), which govern physical systems across various fields, offers substantial benefits for understanding the spatial-temporal evolution of these systems Karniadakis et al. (2021); Wandel et al. (2022); Wei et al. (2024). In recent years, physics-informed neural networks (PINNs) Raissi et al. (2019) have emerged as a new learning paradigm for solving PDEs. PINNs enforce the network to satisfy the constraints given by the PDEs, initial conditions (ICs), and boundary conditions (BCs) via minimizing a composite loss function. The loss in PINNs typically comprises a PDE residual loss evaluated on a sample of scattered interior points (referred to as collocation points), and loss terms for initial and boundary points. In addition, a prominent advantage of PINNs is to seamlessly integrate observational data with physical laws. This integration bridges the gap between data-driven learning and physics-driven learning, thereby offering a powerful framework for a wide range of scientific applications Karniadakis et al. (2021); Shi et al. (2021); Meng et al. (2022); Ashraf et al. (2024); Li et al. (2024); Chu et al. (2024). However, PINNs sometimes have been found to perform poorly and even fail to converge, especially when solving stiff PDEs whose solutions exhibit fast-paced spatial-temporal transitions Krishnapriyan et al. (2021); Kim et al. (2021); Wang et al. (2022b). Previous studies have revealed that a leading cause of failure is attributed to the imbalanced multi-objective loss optimization Wang et al. (2022b, 2021). Typically, the PDE residual loss dominates the total training process and converges faster, while certain condition-fitting losses (collectively referring to the loss terms of IC, BCs, observations, etc.) suffer from vanishing back-propagated gradients Wang et al. (2022b, 2021). The trained model is strongly biased toward yielding a solution with a small PDE residual error, ignoring the accurate fitting of initial and boundary conditions. This is why PINNs often fail to learn the correct solution.
@@
Numerous studies have made efforts to balance the interplay between different terms in the composite loss function by tuning weights for loss components Wang et al. (2021, 2022a); Bai et al. (2023); Dashtbayaz et al. (2024); Song et al. (2024). Wang et al. Wang et al. (2021) propose to dynamically tune weights based on the mean magnitude of the backpropagated gradients of the loss function. This work serves as a foundational contribution, inspiring many follow-up dynamic weighting methods based on different gradient statistics, e.g., standard deviation Maddu et al. (2022), kurtosis Vemuri and Denzler (2023), and Euclidean norms Deguchi and Asai (2023). These gradient statistics-based weighting methods can alleviate the dominance of PDE residual loss by emphasizing the contributions of condition-fitting losses. Specifically, they consider the PDE residual loss as a baseline and assign the weights to condition-fitting losses by relating each condition-fitting loss’s gradients to the PDE loss’s gradients, in order to mitigate the gradient imbalance. However, we notice that these methods are merely based on the pairwise relationship between the PDE residual and each condition-fitting loss. They ignore the intra-relations within condition losses and do not consider the difficulty imbalance for fitting different conditions. This imbalance in condition-fitting difficulty leads to consistently excessive emphasis on easier conditions, thereby neglecting and slowing progress on fitting difficult conditions. Consequently, the model converges slowly and is easily stuck in a local minimum.
@@
To address these issues, we propose a novel Dual-Balanced PINN (DB-PINN) which dynamically adjusts the weights to balance the training process of PINNs. Our proposed DB-PINN integrates inter- and intra-balancing to simultaneously alleviate the imbalance in gradient flow between the governing PDE and enforced conditions at the gradient level as well as the imbalance in the condition-fitting difficulty at the loss-scale level. Specifically, inter-balancing characterizes the relationship between PDE residual loss and all condition-fitting losses based on the gradient statistics of back-propagation to obtain an aggregated weight. On the other hand, intra-balancing aims to evaluate the fitting difficulty of each condition based on its training losses and then allocate the aggregated weight proportionally to these condition-fitting losses. Hence, intra-balancing prevents unnecessary emphasis on fitting easier conditions and allocates larger weights to difficult conditions to accelerate their convergence. Ultimately, the model balances the contributions of PDE residual and condition constraints and also balances the convergence rates among diverse conditions to ensure their losses converge to an identical level.
@@ 2 Background and Related Work
2.0.1 Physics-Informed Neural Networks (PINNs)
@@
2.0.2 Prior Work on Dynamic Weighting Methods for PINNs
@@
The calculating approach directly computes the weights by analyzing the backpropagation gradient dynamics Wang et al. (2021); Maddu et al. (2022) or through the lens of the neural tangent kernel (NTK) theory Wang et al. (2022b). Specifically, Wang et al. Wang et al. (2021) observe that the gradients of PDE residual loss generally attain larger values than the gradients of BC loss and IC loss. Based on such findings, they assign appropriate weights to condition-fitting losses to achieve comparable gradient magnitudes and rewrite the total loss as below:
@@ 3.1 Inter-Balancing Between PDE Residual Loss and Condition-Fitting Losses
Previous studies Wang et al. (2022b); Krishnapriyan et al. (2021); Wang et al. (2021) have revealed that the PDE residual loss generally exhibits dominance over the other loss terms, and this training imbalance severely hinders the training convergence and predictive accuracy of PINNs. To mitigate this imbalanced training, we first employ gradient statistics-based weighting methods to capture the gradient distribution discrepancies between PDE residual loss and every condition loss. Next, the overall gradient discrepancies can be derived by a summation. Thus, the total gradient ratio can be computed and considered as an aggregated weight. Specifically, we formulate the calculation procedure of the aggregated weight by adopting the mean magnitude of gradients (i.e., mean Wang et al. (2021)) as a statistic:
@@ 4.1 PDE Benchmarks
We consider the one-dimensional (1D) wave equation, which conventional PINNs struggle to solve due to its stiffness Wang et al. (2022b); McClenny and Braga-Neto (2020). The wave equation with a boundary condition $u(x,t)=0$ as well as the initial conditions $u(x,0)={\rm sin}(\pi x)+0.5\,{\rm sin}(4\pi x)$ and $u_{t}(x,0)=0$ is formulated as:
@@ 4.3 Main Results
Figure 2 illustrates the training loss and weight curves over training epochs for EW, GW-PINN (std), and DB-PINN (std), revealing the limitations of GW-PINN and highlighting the advantages of DB-PINN. First, compared to EW, GW-PINN enlarges the gap between $\mathcal{L}^{bc}$ and $\mathcal{L}^{ic}$. This is because GW-PINN consistently assigns a larger weight of $\mathcal{L}^{bc}$ than that of $\mathcal{L}^{ic}$, i.e., the curve of $\lambda^{bc}$ is always above $\lambda^{ic}$, even though the model already fits BC better than IC. Second, training losses usually exhibit variances, manifested as sharp fluctuations in the loss curves in Figure 2(a). However, GW-PINN is overly sensitive to instantaneous loss values, and the resulting weights undergo rapid volatility, leading to abrupt spikes. Ultimately, the model fails to reach the optimal. In contrast, DB-PINN effectively mitigates these two issues. In Figure 2(c), DB-PINN avoids overfitting to $\mathcal{L}^{bc}$ and underfitting of $\mathcal{L}^{ic}$; instead, it ensures both condition losses steadily converge at almost the same rate to values below 10e-4. This is attributed to DB-PINN’s capability to dynamically allocate robust weights guided by fitting difficulty through the proposed weight updating strategy. Please refer to the supplementary material for more results.
@@
Table 1 summarizes the results of PINNs using different weighting methods for solving the Klein-Gordon equation. Compared to equal weighting (EW) with the largest errors, other methods significantly reduce prediction errors. This proves the importance of investigating the optimal weights for better balancing several loss components in the PINN optimization. Regardless of the gradient statistics employed, our proposed DB-PINN consistently achieves more stable and accurate predictions than its counterpart GW-PINN. DB-PINN (kurtosis) achieves the least prediction errors. Figure 3(a-c) shows the predictions and errors of EW, UW, and DB-PINN (kurtosis), and it indicates that DB-PINN yields the lowest maximum errors. The snapshots in Figure 3(e-h) prove again that DB-PINN achieves the best prediction performance. Figure 3(d) shows that DB-PINN converges fastest and eventually achieves the least predictive errors.
@@
Table 1 compares the performance of different methods for solving the wave equation. Equal weighting struggles to solve this PDE with the relative $L^{2}$ error around 40%. Uncertainty weighting and SA-PINN achieve a slight improvement in accuracy. GW-PINN enlarges the performance gap with the baseline but exhibits unstable performance related to the statistical metrics. Thus, the choice of gradient statistics plays a significant role in the performance of GW-PINN when solving more intricate stiff PDEs. Fortunately, DB-PINN exhibits a significant advantage in that the performance gaps induced by the employed gradient statistics are greatly reduced. Specifically, DB-PINN consistently obtains the best prediction performance among these different weighting methods. Figure 4(a-c) visually compares the performance of EW, UW, and DB-PINN (kurtosis). DB-PINN achieves lower prediction error than other methods, which is also proved by the comparison in Figure 4(e-h). Figure 4(d) shows that DB-PINN has a fast convergence speed and stable predictive performance with low errors.
@@
Table 1 reports the results for solving the Helmholtz PDE. Compared with GW-PINN using identical statistics, DB-PINN attains lower errors. DB-PINN (kurtosis) yields the least errors with the L2RE of 0.140% and MAE of 0.052%. Hence, the effectiveness of DB-PINN in solving time-independent PDEs also has been validated. Figure 5(a-c) indicates that DB-PINN (kurtosis) has the lowest maximum errors. Figure 5(e-h) shows the snapshots of prediction at the boundary $x=1.00$. DB-PINN (kurtosis) achieves predictions that closely match reference solutions. Figure 5(d) shows that DB-PINN exceeds GW-PINN in both convergence speed and prediction accuracy.
Figure 3: The Klein-Gordon equation. (a-c) The predictions and errors. (d) Prediction error curves. (e-h) Comparison between reference solutions (blue) and predictions (red).
Figure 4: The Wave equation. (a-c) The predictions and errors. (d) Prediction error curves. (e-h) Comparison between reference solutions (blue) and predictions (red).
Figure 5: The Helmholtz equation. (a-c) The predictions and errors. (d) Prediction error curves. (e-h) Comparison between reference solutions (blue) and predictions (red).
@@ 4.4 Ablation Study
We conduct ablation studies on solving the Klein-Gordon equation to evaluate the respective contribution of the dual-balancing and the weight update strategy in DB-PINN. Table 2 reports the results of the ablation study about the dual-balancing mechanism in DB-PINN. Our approach without inter- and intra-balancing is denoted as ‘DB-PINN w/o balancing’, which is equivalent to GW-PINN with the proposed weight update strategy. ‘DB-PINN-Avg’ refers to the DB-PINN without intra-balancing, which allocates the average of the aggregated weight to each condition loss. Regardless of any one of the statistical metrics, ‘DB-PINN w/o balancing’ consistently achieves inferior performance than DB-PINN in Table 1. In particular, ‘DB-PINN-Avg’ obtains poorer accuracy than DB-PINN, which justifies the necessity of balancing the training to learn multiple conditions with different levels of fitting difficulty. Therefore, inter-balancing and intra-balancing are indispensable to addressing two imbalance issues, constituting the dual-balancing mechanism collaboratively.

### v3

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as a new learning paradigm for solving partial differential equations (PDEs) by enforcing the constraints of physical equations, boundary conditions (BCs), and initial conditions (ICs) into the loss function. Despite their successes, vanilla PINNs still suffer from poor accuracy and slow convergence due to the intractable multi-objective optimization issue. In this paper, we propose a novel Dual-Balanced PINN (DB-PINN), which dynamically adjusts loss weights by integrating inter-balancing and intra-balancing to alleviate two imbalance issues in PINNs. Inter-balancing aims to mitigate the gradient imbalance between PDE residual loss and condition-fitting losses by determining an aggregated weight that offsets their gradient distribution discrepancies. Intra-balancing acts on condition-fitting losses to tackle the imbalance in fitting difficulty across diverse conditions. By evaluating the fitting difficulty based on the loss records, intra-balancing can allocate the aggregated weight proportionally to each condition loss according to its fitting difficulty level. We further introduce a robust weight update strategy to prevent abrupt spikes and arithmetic overflow in instantaneous weight values caused by large loss variances, enabling smooth weight updating and stable training. Extensive experiments demonstrate that DB-PINN achieves significantly superior performance than those popular gradient-based weighting methods in terms of convergence speed and prediction accuracy. Our code and supplementary material are available at https://github.com/chenhong-zhou/DualBalanced-PINNs.
@@ 1 Introduction
Accurately and efficiently solving the partial differential equations (PDEs) governing physical systems provides significant advantages for understanding their spatiotemporal evolution Karniadakis et al. (2021); Wandel et al. (2022); Wei et al. (2024). In recent years, physics-informed neural networks (PINNs) Raissi et al. (2019) have emerged as a new learning paradigm for solving PDEs. PINNs enforce the network to satisfy the constraints given by the PDEs, initial conditions (ICs), and boundary conditions (BCs) via minimizing a composite loss function. The loss in PINNs typically comprises a PDE residual loss evaluated on a sample of scattered interior points (referred to as collocation points), and loss terms for initial and boundary points. In addition, a prominent advantage of PINNs is seamless integration of observational data with physical laws. This integration bridges the gap between data-driven learning and physics-driven learning, thereby offering a powerful framework for a wide range of scientific applications Karniadakis et al. (2021); Shi et al. (2021); Meng et al. (2022); Ashraf et al. (2024); Li et al. (2024); Chu et al. (2024). However, PINNs sometimes have been found to perform poorly and even fail to converge, especially when solving stiff PDEs whose solutions exhibit fast-paced spatial-temporal transitions Krishnapriyan et al. (2021); Kim et al. (2021); Wang et al. (2022). Previous studies have revealed that a leading cause of failure is attributed to the imbalanced multi-objective loss optimization Wang et al. (2022, 2021). Typically, the PDE residual loss dominates the total training process and converges faster, while certain condition-fitting losses (collectively referring to the loss terms of IC, BCs, observations, etc.) suffer from vanishing back-propagated gradients Wang et al. (2022, 2021). The trained model is strongly biased toward yielding a solution with a small PDE residual error, ignoring the accurate fitting of initial and boundary conditions. This is why PINNs often fail to learn the correct solution.
@@
Numerous studies have made efforts to balance the interplay between different terms in the composite loss function by tuning weights for loss components Groenendijk et al. (2021); Dashtbayaz et al. (2024); Song et al. (2024). Wang et al. Wang et al. (2021) propose to dynamically tune weights based on the mean magnitude of the backpropagated gradients of the loss function. This work serves as a foundational contribution, inspiring many follow-up dynamic weighting methods based on different gradient statistics, e.g., standard deviation Maddu et al. (2022), kurtosis Vemuri and Denzler (2023), and Euclidean norms Deguchi and Asai (2023). These gradient statistics-based weighting methods can alleviate the dominance of PDE residual loss by emphasizing the contributions of condition-fitting losses. Specifically, they consider the PDE residual loss as a baseline and assign the weights to condition-fitting losses by relating each condition-fitting loss’s gradients to the PDE residual loss’s gradients, in order to mitigate the gradient imbalance. However, we notice that these methods are merely based on the pairwise relationship between the PDE residual and each condition-fitting loss. They ignore the intra-relations within condition losses and do not consider the difficulty imbalance for fitting different conditions. This imbalance in condition-fitting difficulty leads to consistently excessive emphasis on easier conditions, thereby neglecting and slowing progress on fitting difficult conditions. Consequently, the model converges slowly and is easily stuck in a local minimum.
@@
To address these issues, we propose a novel Dual-Balanced PINN (DB-PINN), which dynamically adjusts the weights to balance the training process of PINNs. Our proposed DB-PINN integrates inter- and intra-balancing to simultaneously alleviate the imbalance in gradient flow between the governing PDE and enforced conditions at the gradient level as well as the imbalance in the condition-fitting difficulty at the loss-scale level. Specifically, inter-balancing characterizes the relationship between PDE residual loss and all condition-fitting losses based on the gradient statistics of backpropagation to obtain an aggregated weight. On the other hand, intra-balancing aims to evaluate the fitting difficulty of each condition based on its training losses and then allocate the aggregated weight proportionally to these condition-fitting losses. Hence, intra-balancing prevents unnecessary emphasis on fitting easier conditions and allocates larger weights to difficult conditions to accelerate their convergence. Ultimately, the model balances the contributions of PDE residual and condition constraints and also balances the convergence rates among diverse conditions to ensure that their losses converge to an identical level.
@@ 2 Background and Related Work
Physics-Informed Neural Networks (PINNs)
@@
Prior Work on Dynamic Weighting Methods for PINNs
@@
The calculating approach directly computes the weights by analyzing the backpropagation gradient dynamics Wang et al. (2021); Maddu et al. (2022) or through the lens of the neural tangent kernel (NTK) theory Wang et al. (2022). Specifically, Wang et al. Wang et al. (2021) observe that the gradients of PDE residual loss generally attain larger values than the gradients of BC loss and IC loss. Based on such findings, they assign appropriate weights to condition-fitting losses to achieve comparable gradient magnitudes and rewrite the total loss as follows:
@@ 3.1 Inter-Balancing Between PDE Residual Loss and Condition-Fitting Losses
Previous studies Wang et al. (2022); Krishnapriyan et al. (2021); Wang et al. (2021) have revealed that the PDE residual loss generally exhibits dominance over the other loss terms, and this training imbalance severely hinders the training convergence and predictive accuracy of PINNs. To mitigate this imbalanced training, we first employ gradient statistics-based weighting methods to capture the gradient distribution discrepancies between PDE residual loss and every condition loss. Next, the overall gradient discrepancies can be derived by a summation. Thus, the total gradient ratio can be computed and considered as an aggregated weight. Specifically, we formulate the calculation procedure of the aggregated weight by adopting the mean magnitude of gradients (i.e., mean Wang et al. (2021)) as a statistic:
@@ 4.1 PDE Benchmarks
We consider the one-dimensional (1D) wave equation, which conventional PINNs struggle to solve due to its stiffness Wang et al. (2022); McClenny and Braga-Neto (2020). The wave equation with a boundary condition $u(x,t)=0$ as well as the initial conditions $u(x,0)={\rm sin}(\pi x)+0.5\,{\rm sin}(4\pi x)$ and $u_{t}(x,0)=0$ is formulated as:
@@ 4.3 Main Results
Figure 2 illustrates the training loss and weight curves over training epochs for EW, GW-PINN (std), and DB-PINN (std), revealing the limitations of GW-PINN and highlighting the advantages of DB-PINN. First, compared to EW, GW-PINN enlarges the gap between $\mathcal{L}^{bc}$ and $\mathcal{L}^{ic}$. This is because GW-PINN consistently assigns a larger weight to $\mathcal{L}^{bc}$ than $\mathcal{L}^{ic}$, i.e., the curve of $\lambda^{bc}$ is always above $\lambda^{ic}$, even though the model already fits BC better than IC. Second, training losses usually exhibit variances, manifested as sharp fluctuations in the loss curves in Figure 2(a). However, GW-PINN is overly sensitive to instantaneous loss values, and the resulting weights undergo rapid volatility, leading to abrupt spikes. Ultimately, the model fails to reach the optimal. In contrast, DB-PINN effectively mitigates these two issues. In Figure 2(c), DB-PINN avoids BC overfitting and IC underfitting, and it ensures both condition losses steadily converge at almost the same rate to values below 10e-4. This is attributed to DB-PINN’s capability to dynamically allocate robust weights guided by fitting difficulty with the proposed weight updating strategy. Please refer to the supplementary material for more results.
@@
Table 1 summarizes the results of PINNs using different weighting methods for solving the Klein-Gordon equation. Compared to equal weighting (EW), which has the largest errors, other methods significantly reduce prediction errors. This proves the importance of finding the optimal weights for better balancing several loss components in the PINN optimization. Regardless of the gradient statistics employed, our proposed DB-PINN consistently achieves more stable and accurate predictions than its counterpart, GW-PINN. DB-PINN (kurtosis) achieves the least prediction errors. Figure 3(a-b) shows the point-wise absolute prediction errors of EW and DB-PINN (kurtosis), and it indicates that DB-PINN yields the lower maximum errors. The snapshots in Figure 3(d-f) prove again that DB-PINN achieves the best prediction performance. Figure 3(c) shows that DB-PINN converges fastest and eventually achieves the least predictive errors.
@@
Table 1 compares the performance of different methods for solving the wave equation. Equal weighting struggles to solve this PDE with the relative $L^{2}$ error around 40%. Uncertainty weighting and SA-PINN achieve a slight improvement in accuracy. GW-PINN enlarges the performance gap with the baseline but exhibits unstable performance related to the statistical metrics. Thus, the choice of gradient statistics plays a significant role in the performance of GW-PINN when solving more intricate stiff PDEs. Fortunately, DB-PINN exhibits a significant advantage in that the performance gaps induced by the employed gradient statistics are greatly reduced. Specifically, DB-PINN consistently obtains the best prediction performance among these different weighting methods. Figure 4(a-b) visually compares the performance of EW and DB-PINN (kurtosis). DB-PINN achieves better prediction accuracy, which is also proved by the comparison in Figure 4(d-f). Figure 4(c) shows that DB-PINN has a fast convergence speed and stable predictive performance with low errors.
@@
Table 1 reports the results for solving the Helmholtz PDE. Compared with GW-PINN using identical statistics, DB-PINN attains lower errors. DB-PINN (kurtosis) yields the least errors with the L2RE of 0.140% and MAE of 0.052%. Hence, the effectiveness of DB-PINN in solving time-independent PDEs also has been validated. Figure 5(a-b) indicates that DB-PINN (kurtosis) has the lowest maximum errors. Figure 5(d-f) shows the snapshots of prediction at the boundary $x=1.00$. DB-PINN (kurtosis) achieves predictions that closely match reference solutions. Figure 5(c) shows that DB-PINN exceeds GW-PINN in both convergence speed and prediction accuracy.
Figure 3: The Klein-Gordon equation. (a-b) Point-wise absolute errors. (c) Mean prediction error curves. (d-f) Comparison between reference solutions (blue) and predictions (red).
Figure 4: The Wave equation. (a-b) Point-wise absolute errors. (c) Mean prediction error curves. (d-f) Comparison between reference solutions (blue) and predictions (red).
Figure 5: The Helmholtz equation. (a-b) Point-wise absolute errors. (c) Mean prediction error curves. (d-f) Comparison between reference solutions (blue) and predictions (red).
@@ 4.4 Ablation Study
We conduct ablation studies on solving the Klein-Gordon equation to evaluate the respective contribution of the dual-balancing and the weight update strategy in DB-PINN. Table 2 reports the results of the ablation study about the dual-balancing mechanism in DB-PINN. Our approach without inter- and intra-balancing is denoted as ‘DB-PINN w/o balancing’, which is equivalent to GW-PINN with the proposed weight update strategy. ‘DB-PINN-Avg’ refers to the DB-PINN without intra-balancing, which allocates the average of the aggregated weight to each condition loss. Regardless of any one of the statistical metrics, ‘DB-PINN w/o balancing’ consistently achieves inferior performance than DB-PINN in Table 1. Notice that ‘DB-PINN-Avg’ obtains poorer accuracy than DB-PINN, which justifies the necessity of balancing the training to learn multiple conditions with different levels of fitting difficulty. Therefore, inter-balancing and intra-balancing are indispensable to addressing two imbalance issues, constituting the dual-balancing mechanism collaboratively.